OpenStudy (anonymous):

if your car gets22 miles per gallon how much does it cost to drive 330 miles when gasoline costs 3.00 per gallon

jimthompson5910 (jim_thompson5910):

(22 miles/1 gallon)*( x gallons ) = 330 miles x gallons = (330 miles)*(1 gallon/22 miles) x gallons = (330/22) gallons x gallons = 15 gallons ------------------------------------------- So if you get 22 mpg and you want to drive 330 miles, then you need 15 gallons of gas to do so. (15 gallons)*(3 dollars/1 gallon) = 15*3 = 45 and it will cost you $45 total to fill up those 15 gallons to drive 330 miles. Hopefully this all makes sense. If not, then let me know